the average age of 8 men is increased by 2 years when 2 of them whose ages are 21 years and 23 years are replaced by two new men. the average age of two new men is?
Answer by Theo(13342)

the average of the original 8 men is equal to x

this mans that the sum of their ages divided by the number of men = x

we get S/8 = x

we know that if we subtract (21 and 23) from the sum, and we add the ages of 2 other men, that the average age now becomes (x+2).

formula for this would be:

(S - 44 + y)/8 = (x+2)

y represents the sum of the ages of the 2 new men.

we have 2 equations:

S/8 = x and (S - 44 + y)/8 = (x+2)

if we multiply both sides of each equation by 8, then we get:

S = 8*x and S - 44 + y = 8*(x+2)

Since S = 8*x in the first equation we can susbtitute 8*x for S in the second equation to get:

8*x - 44 + y = 8*(x+2)

We simplify this by removing parentheses to get:

8*x - 44 + y = 8*x + 16

we subtract 8*x from both sides of the equation and we add 44 to both sides of the equation to get:

y = 60

y is the sum of the ages of the 2 new men.

their average age is 60/2 = 30

that should be the answer to your question.